OpenStudy (anonymous):

log_{8} 14^{x-1} =13

OpenStudy (dumbcow):

Rewrite in exponential form \[14^{x-1} = 8^{13}\] take log of both sides \[x-1 = \frac{\log 8^{13}}{\log 14}\]

OpenStudy (anonymous):

and the answer is?

OpenStudy (anonymous):

\[\log_{8}14^{x-1} = 13\]\[(x-1)\log_{8}14 = 13\]\[x - 1 = 13 / \log_{8}14\]\[x = 1 + 13/\log_{8}14\]

OpenStudy (dumbcow):

use your calculator...